zarkam21:

Prove the identity by completing the table below, indicating the steps on the left and the reasoning on the right.

zarkam21:

Vocaloid:

any attempts on the distributive property step yet? A(B+C) = A*B + A*C

zarkam21:

sec(x)*tan(x)+csc(x)*cot(x)

Vocaloid:

close but you have to distribute sec(x)csc(x) to both terms

zarkam21:

sec(x)csc(x)*tan(x)+sec(x)csc(x)*cot(x)

Vocaloid:

good then for the "apply the definitions step" you just re-write everything in terms of sin and cos using the appropriate definitions be careful with parentheses

zarkam21:

confused :/

Vocaloid:

sec(x) = 1/cos(x) csc(x) = 1/sin(x) tan(x) = sin(x)/cos(x) cot(x) = cos(x)/sin(x)

Vocaloid:

so you'd replace everything with its appropriate definition

zarkam21:

okay so sec(x)=1/cos(x)

Vocaloid:

good, keep going you will apply the definitions to everything in the expression sec(x)csc(x)*tan(x)+sec(x)csc(x)*cot(x)

zarkam21:

sec(x) = 1/cos(x) csc(x) = 1/sin(x) tan(x) = sin(x)/cos(x) cot(x) = cos(x)/sin(x)

zarkam21:

like there are no numbers to put right, it will just be sin, cos, etc.

Vocaloid:

yes

zarkam21:

okay so is : sec(x) = 1/cos(x) csc(x) = 1/sin(x) tan(x) = sin(x)/cos(x) cot(x) = cos(x)/sin(x)

zarkam21:

correct?

Vocaloid:

yes, but take the original expression and do the substitutions

zarkam21:

sec(x)csc(x) = 1/cos(x) csc(x) = 1/sin(x) tan(x) = sin(x)/cos(x) cot(x) = cos(x)/sin(x)

zarkam21:

i did the first one, is that correct?

Vocaloid:

sec(x)csc(x) is not equal to 1/cos(x) sec(x) is equal to 1/cos(x) csc(x) is equal to 1/sin(x) therefore sec(x)csc(x) = ?

zarkam21:

sec(x)csc(x) =1/cos(x)+1/sin(x)

Vocaloid:

there's no plus sign it's just (1/cos(x))(1/sin(x))

Vocaloid:

now we have (1/cos(x))(1/sin(x))*tan(x)+sec(x)csc(x)*cot(x) keep going with the rest of the expression

zarkam21:

tan(x) = 1/cos(x)/ 1/sin(x)

Vocaloid:

i guess so but there's a better way to write that, you can just write it as sin(x)/cos(x)

Vocaloid:

anyway now we have (1/cos(x))(1/sin(x))*[sin(x)/cos(x)]+sec(x)csc(x)*cot(x) keep going

zarkam21:

cos(x)/sin(x)

Vocaloid:

good (1/cos(x))(1/sin(x))*[sin(x)/cos(x)]+sec(x)csc(x)*cos(x)/sin(x) keep going with the last sec and csc

zarkam21:

Im having trouble with the sec and csc one

Vocaloid:

sec(x) is equal to 1/cos(x) csc(x) is equal to 1/sin(x) therefore sec(x)csc(x) = ?

zarkam21:

sec(x)csc(x)=1/cos(x) (1/sin(x))

Vocaloid:

good so the entire expression becomes (1/cos(x))(1/sin(x))*[sin(x)/cos(x)]+[1/cos(x)]*[(1/sin(x)]*cos(x)/sin(x)

Vocaloid:

now for the "simplify" step you just cross out expressions when applicable for example, you have (1/sin(x))*[sin(x)] which cancels out to 1

zarkam21:

sin(x)/cos(x)=tan(x)

Vocaloid:

hm, we don't want to re-write anything in terms of cot or tan because that would undo the work we did before

Vocaloid:

look for things you can cancel out between the numerator and denominator

zarkam21:

oh okay..

zarkam21:

then would it be wit cos

Vocaloid:

good, so when you cancel out the sin expression on the left and the cos expression on the right term what do you get?

zarkam21:

the expression I am working with is: (1/cos(x))(1/sin(x))*[sin(x)/cos(x)]+[1/cos(x)]*[(1/sin(x)]*cos(x)/sin(x)

zarkam21:

right

Vocaloid:

yes

zarkam21:

csc^2(x)+sec^2(x)

Vocaloid:

kind of skipping a step there after cancelling out you should get: 1/sin^2(x) + 1/cos^2(x) then the next step asks you to re-write in terms of csc and sec to get csc^2(x)+sec^2(x)

zarkam21:

1/sin^2(x) + 1/cos^2(x) csc^2(x)+sec^2(x)

zarkam21:

so is this it for the simplifying step

Vocaloid:

notice how there's a simplifying step and a step where you apply the definitions of csc and sec

zarkam21:

yes

Vocaloid:

this is the simplifying step 1/sin^2(x) + 1/cos^2(x) this is the step where you apply csc and sec definitions csc^2(x)+sec^2(x)

Vocaloid:

then for the applying pythagorean identities step you just replace csc^2 and sec^2 with their appropriate tan/cot values 

zarkam21:

csc^2(x)=1+cot^2(x) sec^2(x)=tan^2(x)+1

Vocaloid:

good so adding them together = ?

zarkam21:

adding both of the expressions together?

Vocaloid:

yes you have csc^2(x)+sec^2(x) so you must add the expressions together

zarkam21:

1+cot^2(x)+tan^2(x)+1

Vocaloid:

good then for the last simplifying step just combine the like terms (1 and 1)

hotdougstand:

wow alot of steps

zarkam21:

tan^2(x)+cot^2(x)+2

Vocaloid:

good, that's it

zarkam21:

for the pyth identities right

Vocaloid:

1+cot^2(x)+tan^2(x)+1 is the pythagorean identities step tan^2(x)+cot^2(x)+2 is the simplifying step

zarkam21:

thankss!